HUBBLE SPACE TELESCOPE DAILY REPORT #1164 PERIOD COVERED: 0700 EST Daily Status Report as of 172/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOC/96 5330 (High Resolution of Pluto's Albedo Distribution) The Faint Object Camera was used to make observations of the planet Pluto and its moon Charon in two bandpasses. Six images were received and displayed in the OSS. They were reported to be nominal. Some recentering was also reported to have occurred. 1.2 Completed Two Sets of WF/PC-2 5562 (WF/PC-2 Cycle 4 Calibration: Internal Monitor 3: Darks) The WF/PC-2 was used to obtain several 30-minute long, dark exposures to provide ongoing calibration of the CCD dark current and to monitor and characterize the evolution of hot pixels. Five images were received and displayed in the OSS. The images were reported to look okay, except one observation was reported to have a residual diffraction spike from a bright star. 1.3 Completed WF/PC-2 5655 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 2: Flats Part 2) The WF/PC-2 was used to make dark flat calibration observations. Two images were received and displayed. Both images were reported to be nominal. 1.4 Completed Three Sets of WF/PC-2 5446 (The Nuclei of Nearby S0 and Spiral Galaxies) The WF/PC-2 was used to make observations of galaxies NGC-7090, NGC-247, NGC-3056, and NGC-2997. The observations were executed as scheduled. Six images were received and displayed in the OSS. They all were reported to be nominal. More observations are still waiting to be processed. No anomalies were noted. 1.5 Completed WF/PC-2 5479 (Sub-Arcsecond Structure in AGNs) The WF/PC-2 was used to make subarcsecond observations of the active galactic nuclei PKS0518-45. The baseline guide star acquisition failed to acquire the guide star in Fine Lock. The guide star was acquired in course track. ( See observations note 2.1) The observation did not show an obvious target. 1.6 Completed WF/PC-2 5450 (Host Galaxies and Cluster Environments of Quasars) The WF/PC-2 was used to make observations of Quasar QSO-235058 to study the properties of its host galaxy and surrounding environment. The observations were executed as scheduled. Seven observations were received and displayed. They were reported to be nominal. Some recentering was also reported to have occurred. 1.7 Completed WF/PC-2 5560 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 1) The WF/PC-2 was used to make several internal bias, K-Spot, and flat calibration observations to monitor the CCD camera electronics, the optical alignment, and to check for contamination. A total of eight images was taken. They all were reported to be nominal 1.8 Completed WF/PC-2 5427 (Calibration of Nearby Galaxy Type IA Supernovae as Standard Candles) The WF/PC-2 was used to make observations of nearby galaxy NC-4496-A-F as part of an effort to find Cepheids in this galaxy which has produced a Type I supernovae. The observations were executed as scheduled, and still remain on Science Tape Recorder #1. They will be analyzed later. No anomalies were noted. 1.9 Completed FOS/BL 5530 (FOS Location of Spectra and Aperture) The Faint Object Spectrograph (Blue Side) was used to make Y-base maps of the FOS detectors using internal lamps. The purpose is to determine the location of the spectra on the diodes. The data still remain on Science Tape Recorder #1. It will be analyzed later. No anomalies were noted. 1.10 Complete GHRS/2 5596 (GHRS Side 2 Sensitivity Calibration (1st Order Gratings): Cycle 4) The High Resolution Spectrograph (Side Two) was used to measure its sensitivity to provide information on temporal variations in throughput since the installation of COSTAR. Observations were made of the standard test object Mu-Columbae. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 6 At 171/2210Z the FGS Fine Lock acquisition failed to Coarse Track due to the step size limit being exceeded. (See Note 1.5) Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled :22 Successful :22 2.3 Significant HSTARs: None 2.4 Operations Notes: Another quiet operational day was experienced as each HST subsystem reported no anomalous activity. 3.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-